28. test A population proportion is .40. A random sample of size 200 will be taken and the sa proportion p will be used to estimate the population proportion. a. What is the probability that the sample proportion will be within +.03 of the popu tion proportion? b. What is the probability that the sample proportion will be within +.05 of the popula- tion proportion? Aceume that the population proportion is .55. Compute the standard error of the proportion. 10
Answer #1

28)\

a)

for normal distribution z score =(phat-p)/\sigmap
here population proportion=     p= 0.400
sample size       =n= 200
std error of proportion=\sigmap=\sqrt{}(p*(1-p)/n)= 0.0346

hence probability that sample proportion will be within -/+ 0.03 of population proportion:

probability = P(0.37<X<0.43) = P(-0.87<Z<0.87)= 0.8078-0.1922= 0.6156

b)

probability that sample proportion will be within -/+ 0.05 of population proportion:

probability = P(0.35<X<0.45) = P(-1.44<Z<1.44)= 0.9251-0.0749= 0.8502